OpenStudy (anonymous):

Complete the 2 column proof. Given: x over 6+2=15 Prove: x=78

OpenStudy (anonymous):

what is this???

OpenStudy (anonymous):

What do you mean?

OpenStudy (anonymous):

the problem is x/6+2=15 => x/6=15-2=13 => x=13*6=78

OpenStudy (anonymous):